OpenStudy (anonymous):

1/3, 5/12, 3/8 in order from least to greatest

OpenStudy (anonymous):

1/3, 3/8,5/12

OpenStudy (anonymous):

convert to fractions with LCM = 24 1/3 = 8/24 5/12 = 10/24 3/8 = 9/24 so the order is now obvious

OpenStudy (anonymous):

Okay, so 1/3 is 8/8 * 1/3 as multiplying anything over itself does not change the fraction. So 1/3 = 8/24. 3/8 * 3/3 = 9/24 5/12 * 2/2 = 10/24 So if we line these up, we have 8/24 (1/3), 9/24 (3/8), 10/24 (5/12). Hope this helps. If it does, please choose good answer. Visit my website for free math videos. http://www.tutorsean.net

OpenStudy (anonymous):

hope these have helped